OpenStudy (anonymous):

write an equation in slope-intercept form for a line that is perpendicular to the given line and that passes through the given point

OpenStudy (anonymous):

the problems are (-10,3); y= -2x-8 (12,8);y=-3x-9 (-1,7);y=-1/4x+6 (4,-9);-x + 2y=10 (2,1);-2x +3y=-12 (-2,3);y=-5x +1

OpenStudy (anonymous):

refresh my memory what is the slope-intercept formula?

OpenStudy (anonymous):

y-mx+b

OpenStudy (anonymous):

y=mx+b*

OpenStudy (anonymous):

you know what all the positions are so to find slope u need another point so i am only going to explain ONE and u will do the rest ok?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

(-10,3); y= -2x-8 sub in the -10 for the x position wat do ya get?? when you solve it

OpenStudy (anonymous):

you get y=20-8...y=12

OpenStudy (anonymous):

so now u have your other point and u take (-10,12) and put it in the slope formula to find the slope.. y2 - y1 divided by x2 -x1 so do that

OpenStudy (anonymous):

i don't get that part

OpenStudy (anonymous):

can you like do the next problem with steps and i will do the other ones by looking at your steps

OpenStudy (anonymous):

oh sorry i have to go i cant

OpenStudy (anonymous):

:( ok

OpenStudy (anonymous):

To make a perpendicular line, you just take the negative reciprocals of the slope